OpenStudy (anonymous):

How to solve for z in terms of x and y? 2(2)(x^2+y^2)(2x+2yz)=100[xz+y]

OpenStudy (ajprincess):

\(2(2)(x^2+y^2)(2x+2yz)=100(xz+y)\) Is this the question?

OpenStudy (anonymous):

yup.

OpenStudy (ajprincess):

\(4(2x^3+2xy^2+2x^2yz+2y^3z)=100(xz+y)\) \(2x^3+2xy^2+2x^2yz+2y^3z=25xz+25y\) \(2x^3+2xy^2-25y=25xz-2x^2yz-2y^3z\) Factorise RHS. Take z as the common factor from \(25xz-2x^2yz-2y^3z\) Divide by \(25x-2x^2y-2y^3\) both sides.

hartnn (hartnn):

or what u can do is,just for convenience, take \[a=4(x^2+y^2)\] to get \[2ax+2ayz=100xz+100y\] solving for z here is easier in the end u can resubstitute the value of a.....

OpenStudy (anonymous):

Thank you guys! I get it now:)

OpenStudy (ajprincess):

yw.:)